During another safety crash test, a 2550 kg truck traveling east at 19.5 m/s crashes into a 1780 kg compact car that is traveling west at 39.2 m/s. The two crumple together and travel as one unit. Find their velocity and direction after impact.
1
Expert's answer
2020-11-02T09:24:59-0500

Let's choose eastward as the positive direction. Then, we can find the velocity of the crumple after the impact from the Law of Conservation of Momentum:


"m_1v_1-m_2v_2=(m_1+m_2)v,"

here, "m_1=2550\ kg" is the mass of the truck, "v_1=19.5\ \dfrac{m}{s}" is the velocity of the truck, "m_2=1780\ kg" is the mass of the car, "v_2=39.2\ \dfrac{m}{s}" is the velocity of the car, "v" is the velocity of the crumple after the impact.

Then, from this equation we can find "v":


"v=\dfrac{m_1v_1-m_2v_2}{m_1+m_2},""v=\dfrac{2550\ kg\cdot 19.5\ \dfrac{m}{s}-1780\ kg\cdot 39.2\ \dfrac{m}{s}}{2550+1780}=-4.63\ \dfrac{m}{s}."

The sign minus means that the crumple after the impact traveling west.

Answer:

"v=4.63\ \dfrac{m}{s}," west.